OpenStudy (anonymous):

How would you solve x for 2bx + 3ax = b^2 - a^2?

OpenStudy (anonymous):

OOOH OUCH...I would do this: \[\eqalign{ &2bx+3ax=b^2-a^2 \\ &(2b+3a)x=b^2-a^2 \\ &x=\frac{b^2-a^2}{2b+3a} }\]

OpenStudy (anonymous):

Yep got it thanks @KeithAfasCalcLover

OpenStudy (anonymous):

Anytime @AnImEfReaK !

OpenStudy (anonymous):

I don't believe you can actually solve for it for real. Its actually a literal equation which happens to take the form of what is called an infinite elliptic cone. However, that's not important here. The only thing I really know to do here is move everything to one side first. So, a^2-b^2+2bx+3ax=0 Then, we can sort of factor by grouping. It will look like this. (a^2-b^2)+(2bx+3ax)=0 (a+b)(a-b)+x(2b+3a)=0 Now, move (a+b)(a-b) to the right hand side x(2b+3a)=-(a+b)(a-b) Divide by (2b+3a). x=-(a+b)(a-b)/(2b+3a)

OpenStudy (anonymous):

Basically the same explanation as KeithAfasCalcLover only looking at it from a longer route

OpenStudy (anonymous):

Yeah expansion.

OpenStudy (anonymous):

But for the simplest form, you would always multiply out the numerator before calling it finished.